OpenStudy (anonymous):

three forces with magnitudes 100, 50, and 80 pounds, act on an object an angles of 50, 160, and -20 degrees respectively

OpenStudy (dumbcow):

Net force on object is dictated by sum of horizontal and vertical components \[F = <F_x, F_y>\] \[F_x = 100\cos(50)+50\cos(160)+80\cos(-20)\] \[F_y = 100\sin(50)+50\sin(160)+80\sin(-20)\]

OpenStudy (anonymous):

thank you so much! that makes a lot of sense! thanks again :)

OpenStudy (dumbcow):

yw